please help me on this: the complement of an angle is 18 degree more than 3 times the measure of the angle. find the measures of both angles.
Answer by stanbon(75887)

the complement of an angle is 18 degree more than 3 times the measure of the angle. find the measures of both angles.
=====
angle:: x
complement:: 90-x
------
Equation:
90-x = 3x+18
------
4x = 72
x = 18 (measure of the angle)
90-x = 72 (measure of the complement)
----------------
Checking:::
3x+18 = 3*18+18 = 54+18 = 72
-----------
Cheers,
Stan H.
------------